Accessibility Assistance

Skip to Content

Information Technology
George A. Smathers Libraries
University of Florida
P.O. Box 117000
Gainesville, FL 32611-7000

Telephone: (352) 273-2800
Fax: (352) 392-2354

Information Technology Staff Meeting Notes
November 1, 2011

Present: Shruti Bagri, Thomas Bielicke, Cynthia Bowker, Laurie Brennan, Will Chaney, Logan Clapp, Debra Fetzer, Tom Minton, Chris Nicolich, Amy Polk, Mark Sullivan, Evan Wack
  1. Monthly IT update - cancelled due to Smathers 1A room maintenance.
  2. Open Enrollment - ongoing, November 7-18.
  3. Funding requests for MSL netbooks approved.
  4. November 17, Judith Roberts will be out to ECOB IT to discuss the training needs of and to get familiar with our department.
  5. New campus web site is up, highlighting the redesign of the UF web page:
Ongoing IT Updates

  1. App Streaming - testing DLC software; working great. We're also testing Cataloging software.
  2. Migration - new cart delivered for West netbooks. Under configuration. UAC turned off on all staff machines. West Circulation, Cataloging and Administration are next on the migration list.
  3. 1A refresh - wall behind the podium is totally gutted.
  4. VIVO equipment - to be reallocated.
  5. VDI Project - thin clients set up in ECOB and West for testing. Issue with log on time.
  6. Images - removing language packs from all public images. Adding .Net framework to staff image. Reboot testing ongoing.
  7. Smart Control - MSL laptop, West netbook issues - gets two Mac addresses, wired network cannot turn off. Run as service?
  8. Privilege Authority - testing.
  1. Document disposition database - new requirements from Judy Russell - should take about two weeks.
  2. Netbook NIC enabler - Chris is working on it.
  3. DLC - working with FCLA.
  4. Film Tracker - will be a replacement for two old systems. The second portion of it will go to Shruti.
  5. Storage Gap Tracker - working on requirements and database design.
  6. ARL Stats - done.
  7. ARES-Sakai authentication - Michael Jay is working on it.
Network/Server Administration
  1. People Counter - installation almost done.
  2. New server for OLE project: It is up and running.
  3. New departmental VPN will be deployed for Library IT use. All server access will be shifted to require the departmental VPN.
  4. VDI/VMware pilot project - configured test linked clone pool last week.
  5. ALADN website/list move - still in progress.
  6. Local GitHub install for code management: - configuring it now.
  7. Blogs moved to new server.
  8. Red Hat servers now being deployed as preferred distro for all Library Linux servers. All patch management now done remotely through satellite server at CNS.